10 years ago the father was 7times older than his son.2 years hence 2 times the fathers age will be 5 times the sons age.what are there present age.
Answer by ewatrrr(24785)


Hi,
Let x and 7x represent the son's and father's ages 10 year ago.
Question states***
2(7x+12) = 5(x + 12)
solving for x
14x + 24 = 5x = 60
9x = 36
x = 4, son's age 10 years ago. His father 28, ten year ago
Son is NOW 14 and his father 38
CHECKING our Answer using status 2 years from NOW
2*(38+2) = 5*(14+2) = 80